OpenStudy (anonymous):

Anyone willing to help solve Matrices? Solve each matrix equation 15.) 3[2 0] -2X=[-10 5] [-1 5] [0 17] 16.) 4X+ [1 3] = [-3 11] [-7 9] [5 -7]

OpenStudy (anonymous):

Is the first one [6 0] [-1 5]? well for the first part?

OpenStudy (anonymous):

\[3 \left[\begin{matrix}2 & 0 \\ -1 & 5\end{matrix}\right]=\left[\begin{matrix}6 & 0 \\ -3 & 15\end{matrix}\right]\]

OpenStudy (anonymous):

Ohh. . .Don't know how you got that, but ugh, thanks.

OpenStudy (anonymous):

\[\left[\begin{matrix}6 & 0 \\ -3 & 15\end{matrix}\right]-\left[\begin{matrix}-10 & 5 \\ 0 & 17\end{matrix}\right]=2 X\]

OpenStudy (anonymous):

\[2 X=\left[\begin{matrix}6+10 & 0-5 \\ -3-0 & 15-17\end{matrix}\right]\]

OpenStudy (anonymous):

\[2 X=\left[\begin{matrix}16 & -5 \\ -3 & -2\end{matrix}\right]\]

OpenStudy (anonymous):

\[X=\left[\begin{matrix}8 & -5/2 \\ -3/2 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

Thanks for your patience. <3

OpenStudy (anonymous):

yw